Write an equation of the line through the given points.


26. (0,2),(3,8)
Any help with the given equation would Greatly appreciated.

Answer by checkley71(8403)

FIRST WE FIND THE SLOPE
(Y2-Y1)/X2-X1)=(8-2)/(3-0)=6/3=2 ANSWER FOR THE SLOPE (m).
NOW SUBSTITUTE ONE SET OF X,Y COORDINATES & SOLVE FOR THE Y INTERCEPT (b).
Y=mX+b
2=2*0+b
2=b ANSWER FOR THE Y INTERCEPT.
THUS THE LINE EQUATION IS:
Y=2X+2